# Neural Collaborative Filtering to Detect Anomalies in Human Semantic Trajectories

### v2

@@ Abstract.
Human trajectory anomaly detection has become increasingly important across a wide range of applications, including security surveillance and public health. However, existing trajectory anomaly detection methods are primarily focused on vehicle-level traffic, while human-level trajectory anomaly detection remains under-explored. Since human trajectory data is often very sparse, machine learning methods have become the preferred approach for identifying complex patterns. However, concerns regarding potential biases and the robustness of these models have intensified the demand for more transparent and explainable alternatives. In response to these challenges, our research focuses on developing a lightweight anomaly detection model specifically designed to detect anomalies in human trajectories. We propose a Neural Collaborative Filtering approach to model and predict normal mobility. Our method is designed to model users’ daily patterns of life without requiring prior knowledge, thereby enhancing performance in scenarios where data is sparse or incomplete, such as in cold start situations. Our algorithm consists of two main modules. The first is the collaborative filtering module, which applies collaborative filtering to model normal mobility of individual humans to places of interest. The second is the neural module, responsible for interpreting the complex spatio-temporal relationships inherent in human trajectory data. To validate our approach, we conducted extensive experiments using simulated and real-world datasets comparing to numerous state-of-the-art trajectory anomaly detection approaches.
Recommendation Systems, Collaborative Filtering, Neural Networks, Deep Learning
@@ 1. Introduction
    We perform extensive experiments on three highly sparse and extremely imbalanced datasets  (Amiri et al., 2023, 2024; Zheng et al., 2008a) to demonstrate the effectiveness and the promise of our neural collaborative filtering approaches on the user-level spatial-temporal anomaly detection domain.
@@ 2.1. Trajectory Anomaly Detection
anomalyOver the past few decades, a broad spectrum of trajectory outlier detection methstatistical, knowledge-driven,. In this section, we approachegorize these approaches into conventional techniques and machine learning-driven methodologies.
@@ 4. Methodology
Figure 5. The illustration of our proposed NCF model. (Left) Extraction of spatio-temporal embeddings from input trajectories; (Middle) Neural Collaborative Filtering with two-level of prediction; (Right) Anomaly scores based on weighted two-level of prediction.
@@ 4.1. Neural Collaborative Filtering Model
In our Spatial-Temporal model, we extend beyond the conventional method by integrating not only user and POIs latent variables but also additional latent features specifically designed to capture spatial and temporal dynamics. This includes geographic information that captures the spatial context of the interactions, as well as temporal attributes that detail the duration of stays and specific visit times. The geographic information provides a framework for understanding the spatial relationships and locations of the staypoints of interest, while the temporal attributes offer insights into the timing and frequency of visits. After the initial embedding layer, the model utilizes multiple fully connected layers, which function as neural collaborative filtering layers.
The overall MLP module was trained on the training dataset $\mathcal{T}^{Train}$, with the objective of capturing the representation of normal patterns. Once trained, any deviations from these established patterns are indicative of anomalies in the test dataset $\mathcal{T}^{Test}$. In the following section, we describe how to measure these deviations and to identify anomalous users.
@@ 5. Experimental Results and Environmental setup
In this section, we conduct a comprehensive evaluation of our neural collaborative filtering model using multiple simulated and real-world datasets. The subsequent parts of this section will provide detailed information about each dataset, including their anomaly characteristics and the specific preprocessing steps applied. Additionally, we will outline the anomalous settings and configurations in the simulated datasets used for our model.
@@ 5.1.1. Agent-Based Simulation of Patterns of Life
The Patterns of Life simulation is an existing simulation that was designed to emulate human needs and behaviors in an urban environment (Züfle et al., 2023; Kohn et al., 2023). Within the simulation, agents engage in activities that mimics real-life actions, such as attending work, forming friendships, and participating in social gatherings. These agents navigate a virtual environment modeled on real-world settings like roads and buildings, sourced from OpenStreetMap.

### v3

@@ Abstract.
Human trajectory anomaly detection is critical for applications such as security surveillance and public health, yet most existing methods focus on vehicle-level traffic, with limited attention to human-level trajectories. Due to the inherent sparsity of human trajectory data, machine learning approaches are favored for detecting complex patterns. However, concerns about model biases and robustness have highlighted the need for more transparent and explainable solutions. In this paper, we propose a lightweight anomaly detection model specifically designed to detect anomalies in human trajectories. We propose a Neural Collaborative Filtering approach to model and predict normal mobility. Our method is designed to model users’ daily patterns of life without requiring prior knowledge, thereby enhancing performance in scenarios where data is sparse or incomplete, such as in cold start situations. Our algorithm consists of two main modules. The first is the collaborative filtering module, which applies collaborative filtering to model normal mobility of individual humans to places of interest. The second is the neural module, responsible for interpreting the complex spatio-temporal relationships inherent in human trajectory data. To validate our approach, we conducted extensive experiments using simulated and real-world datasets comparing to numerous state-of-the-art trajectory anomaly detection approaches.
Recommendation Systems, Collaborative Filtering, Neural Networks
@@ 1. Introduction
    We perform extensive experiments on three highly sparse and extremely imbalanced datasets  (Amiri et al., 2023, 2024a; Zheng et al., 2008a) to demonstrate the effectiveness and the promise of our neural collaborative filtering approaches on the user-level spatial-temporal anomaly detection domain.
@@ 2.1. Trajectory Anomaly Detection
Over the past few decades, a broad spectrum of trajectory anomaly detection methods has been introduced. In this section, we systematically categorize these approaches into statistical, knowledge-driven, and machine learning approaches.
@@ 4. Methodology
Figure 5. The illustration of our proposed NCF model. (Left) Extraction of spatio-temporal embeddings from input trajectories; (Middle) Neural Collaborative Filtering with two-level of prediction; (Right) Anomaly scores based on weighted prediction.
@@ 4.1. Neural Collaborative Filtering Model
In our Spatial-Temporal model, we extend beyond the conventional method by integrating not only user and POIs latent variables but also additional latent features specifically designed to capture spatial and temporal dynamics. This includes geographic information that captures the spatial context of the interactions, as well as temporal attributes that detail the duration of stays and specific visit times. The geographic information provides a framework for understanding the spatial relationships and locations of the staypoints of interest, while the temporal attributes offer insights into the timing and frequency of visits. After the initial embedding layer, the model utilizes multiple fully connected layers, which function as neural collaborative filtering layers. The overall MLP module was trained on the training dataset $\mathcal{T}^{Train}$, with the objective of capturing the representation of normal patterns. Once trained, any deviations from these established patterns are indicative of anomalies in the test dataset $\mathcal{T}^{Test}$. In the following section, we describe how to measure these deviations and to identify anomalous users.
@@ 5. Experimental Results and Environmental setup
In this section, we conduct a comprehensive evaluation of our neural collaborative filtering model using multiple simulated and real-world datasets. The subsequent parts of this section will provide detailed information about each dataset, including their anomaly characteristics and the specific preprocessing steps applied. Additionally, we will outline the anomalous settings and configurations in the simulated datasets used for our model. resent a thorough evaluation of our model’s performance.
@@ 5.1.1. Agent-Based Simulation of Patterns of Life
The Patterns of Life simulation is an existing simulation that was designed to emulate human needs and behaviors in an urban environment (Züfle et al., 2023; Kohn et al., 2023; Amiri et al., 2024b). Within the simulation, agents engage in activities that mimics real-life actions, such as attending work, forming friendships, and participating in social gatherings. These agents navigate a virtual environment modeled on real-world settings like roads and buildings, sourced from OpenStreetMap.